# Beam energy dependence of transverse momentum distribution and elliptic flow in Au-Au collisions using HYDJET++ model

### v1

@@ III.3 Elliptic flow ($v_{2}$)
The heavier quarks like charm and beauty interact with the QGP medium and experience a drag, which affects their transport coefficients. This leads to a mass-dependent pertinent elliptic flow of the heavy quarks at the end of the QGP phase [42, 43]. A similar phenomenon is possible for the strange quarks at BES energies due to the lower temperatures of these systems. Hence, in this work, we have treated the strange and non-strange hadrons differently to check the impact of quark content on the scaling factor ”k”.
@@
Figure 17: Relation between scaling factor k and beam energy. Lines show the fitting function from eq.  (15).
Figure 18: $v_{2}$ as a function of $p_{T}$ for $\sqrt{s_{NN}}=$130 GeV.
The elliptic flow($v_{2}$) at lower $p_{T}$($p_{T}<1.5$ GeV/c) shows a dependence on the mass of the particle species. The particles with lower mass have a larger value of $v_{2}$ i.e. ($v_{2}(\pi^{+})>v_{2}(k^{+})>v_{2}(p)$). A similar trend is also observed for corresponding anti-particles [45, 46]. A mass ordering in $v_{2}$ has been observed in the Au-Au collision at 200 GeV [47, 48]. Recent experiments like RHIC BES have also observed a mass ordering in $v_{2}$ at lower collision energies [18]. To validate our approach of treating strange and non-strange hadrons separately, we should check for mass ordering in our calculations. In Fig. 12 to Fig. 16 we have shown $v_{2}$ as a function of $p_{T}$ for $\sqrt{s_{NN}}$= 62.4, 39.0, 27.0, 19.6, and 11.5 GeV. The $v_{2}$ of all the particles are plotted together for a clear comparison. The $v_{2}$ of particles are represented by solid markers and the $v_{2}$ of antiparticles are shown by open markers to get an idea about the difference in $v_{2}$ of particles and antiparticles, more details can be found in the legend. A clear mass ordering can be seen in the HYDJET++ calculation in all five collision energies. Mass ordering is observed up to $p_{T}$ = 1.5 GeV/c. After 1.5 GeV/c the $v_{2}$ of protons increases rapidly and surpasses the $v_{2}$ of pions and kaons. At about 2 GeV/c an opposite trend is observed i.e ($v_{2}(\pi^{+})<v_{2}(k^{+})<v_{2}(p)$). For $\pi^{\pm}$ and $k^{\pm}$ the $v_{2}$ is the same for both particles and anti-particles, while at 11.5 GeV $k^{-}$ has a higher $v_{2}$ than $k^{+}$. Protons have relatively larger $v_{2}$ than the anti-protons at all collision energies.
We have found that the value of scaling factor $k$ in minimum bias collisions has the following relation with beam energy.
@@
The value of $a$ is almost the same for all particles, while the value of $b$ for non-strange hadrons is greater than the strange hadrons. It indicates that in the HYDJET++ framework, the magnitude of azimuthal spatial and momentum anisotropy (which determines the elliptic flow($v_{2}$)) is dependent on quark content, which suggests that the strange quarks also experience a drag inside the QGP similar to the charm and beauty quarks. The collective behavior is weaker for the lower energies of the Beam Energy Scan range. Hence, the particles are emitted uniformly in all directions [49] which reduces their spatial and momentum anisotropy which reflects as a reduced ”k” value [50]. A clear depiction can be found in Fig. 17.
Our analysis of $v_{2}$ shows a trend in variation of $k$ with beam energy as discussed earlier. We have studied the top RHIC energy of 200 GeV and the lower energies. It is essential to verify this trend at intermediate collision energy. We have chosen 130 GeV as the intermediate energy. In Fig. 18 we have shown $v_{2}$ as a function of $p_{T}$ for minimum bias collisions (0-80%) at $\sqrt{s_{NN}}$ = 130 GeV. The input value of $k$ is taken according to eq.  (15), and $\delta$ and $\epsilon$ are calculated accordingly. The baryonic chemical potential and chemical freezeout temperatures are taken according to eq.  (3). The rest input parameters are same as default values for $\sqrt{s_{NN}}=200$ GeV. The combined elliptic flow of non-strange hadrons ($\pi^{\pm}$ and $p,\bar{p}$) are represented with red markers, and the combined $v_{2}$ of $k^{\pm}$ is shown with blue markers. The $v_{2}$ of $\pi^{\pm}$ and $p,\bar{p}$ is higher than the combined flow of charged kaons ($k^{\pm}$). This is due to the higher value of $\epsilon$ for non-strange hadrons than strange ones which leads to a higher combined $v_{2}$. The results are compared with the combined $v_{2}$ of charged hadrons measured experimentally by the STAR detector [51].
@@ IV Summary
We have performed a comprehensive study of bulk properties of medium like $p_{T}$ spectra and elliptic flow produced in Au-Au collisions at different beam energies. HYDJET++ model shows a good match with the experimentally observed particle ratios with suitable chemical potentials and chemical freezeout temperatures. HYDJET++ calculations of $p_{T}$ spectra show a good match with experimental data for central and semi-central collisions. The model slightly under-predicts $p_{T}$ spectra of $p$ and $\bar{p}$. The contribution of hard and soft processes decreases with decreasing beam energy. The contributions from the hard part decrease more rapidly than the soft part, at lower energies like 11.5 GeV $p_{T}$ spectra is entirely determined by the soft processes. The contribution of the soft part in the invariant yield of protons shows an opposite trend, it increases with decreasing beam energy.
HYDJET++ results on elliptic flow show a good match with experimental data. The model calculations as well as the experimental show a beam energy dependence of $v_{2}$ after 1 GeV/c, the $v_{2}$ in these regions increases with beam energy. No beam energy dependence is seen for $p_{T}<1$ GeV/c. In the regime of low $p_{T}$, a mass ordering in $v_{2}$ is also observed in our model calculations i.e. ($v_{2}(\pi^{+})>v_{2}(k^{+})>v_{2}(p)$), this trend is also seen for the anti-particles. The mass ordering is not observed after 1.5 GeV/c. The difference between $v_{2}$ of particles and anti-particles is not observed for mesons, while a slight difference in $v_{2}$ is seen for $p$ and $\bar{p}$ ($v_{2}(p)>v_{2}(\bar{p})$). The azimuthal spatial anisotropy increases with collision energy as shown in eq. (15).The azimuthal spatial anisotropy ($\epsilon$) at a given collision energy is the same for $\pi^{\pm},p$ and $\bar{p}$, while the charged kaons have a lesser value of $\epsilon$, which suggests that the relatively heavier strange quarks interact with the QGP medium and experience a drag that affects the final spatial and momentum anisotropy of strange hadrons.

### v2

@@ III.3 Elliptic flow ($v_{2}$)
The strange quarks can have a lower collectivity than up/down quarks due to their higher mass. The flow of the quarks can’t be directly observed in experiments. To test this aspect, we used the scaling parameter “$k$” as a proxy for collective behavior. Hence, in this work, we have treated the strange and non-strange hadrons differently to check the impact of quark content on the scaling factor “$k$”.
@@
The elliptic flow($v_{2}$) at lower $p_{T}$($p_{T}<1.5$ GeV/c) shows a dependence on the mass of the particle species. The particles with lower mass have a larger value of $v_{2}$, i.e. ($v_{2}(\pi^{+})>v_{2}(k^{+})>v_{2}(p)$). A similar trend is also observed for corresponding anti-particles [44, 45]. A mass ordering in $v_{2}$ has been observed in the Au-Au collision at 200 GeV [46, 47]. Recent experiments like RHIC BES have also observed a mass ordering in $v_{2}$ at lower collision energies [18]. To validate our approach of treating strange and non-strange hadrons separately, we should check for mass ordering in our calculations. In Fig. 12 to Fig. 16, we have shown $v_{2}$ as a function of $p_{T}$ for $\sqrt{s_{NN}}$= 62.4, 39.0, 27.0, 19.6, and 11.5 GeV. The $v_{2}$ of all the particles are plotted together for a clear comparison. The $v_{2}$ of particles are represented by solid markers, and the $v_{2}$ of antiparticles are shown by open markers to get an idea about the difference in $v_{2}$ of particles and antiparticles, more details can be found in the legend. A clear mass ordering can be seen in the HYDJET++ calculation in all five collision energies. Mass ordering is observed up to $p_{T}$ = 1.5 GeV/c. After 1.5 GeV/c the $v_{2}$ of protons increases rapidly and surpasses the $v_{2}$ of pions and kaons. At about 2 GeV/c an opposite trend is observed i.e ($v_{2}(\pi^{+})<v_{2}(k^{+})<v_{2}(p)$). For $\pi^{\pm}$ and $k^{\pm}$ the $v_{2}$ is the same for both particles and anti-particles, while at 11.5 GeV $k^{-}$ has a higher $v_{2}$ than $k^{+}$. Protons have a relatively larger $v_{2}$ than the anti-protons at all collision energies.
We have found that for low $p_{T}$ ($p_{T}<1.5$ GeV), the value of the scaling factor “$k$” in minimum bias collisions has the following relation with beam energy.
@@
The value of $a$ is almost the same for all particles, while the value of $b$ for non-strange hadrons is greater than the strange hadrons. It indicates that in the HYDJET++ framework, the magnitude of azimuthal spatial and momentum anisotropy (which determines the elliptic flow($v_{2}$)) is dependent on quark content. A lower $k$ for strange hadrons suggests a lower $v_{2}$ of strange quarks before hadronization. A similar observation was made by J. Song $et$ $al$ [48]. The study extracted the $v_{2}$ of quarks based on the Equal Velocity quark Combination (EVC) mechanism and found that strange quarks have smaller $v_{2}$ compared to up/down quarks. However, more studies are required to know the exact reason for such behavior. The collective behavior is weaker for the lower energies of the Beam Energy Scan range. Hence, the particles are emitted uniformly in all directions [49], which reduces their spatial and momentum anisotropy, which reflects as a reduced “$k$” value [50]. A clear depiction can be found in Fig. 17.
In Fig. 18, we have shown the elliptic flow of $k^{+}$, $\pi^{+}$, and p in minimum bias Au-Au collisions at 62.4 GeV with different values of “$k$”. It can be seen that the HYDJET++ calculations are sensitive to “$k$”, a small change in “$k$” can have a notable change in $v_{2}$. The effect is more distinguished for the mesons than protons. If the value of “$k$”, which was used for non-strange hadrons is used for strange hadrons, it would significantly overpredict the data, and vice versa. Hence, different values of $k$ should be used for both cases to better describe the data.
@@ IV Summary
We have performed a comprehensive study of bulk properties of medium like $p_{T}$ spectra and elliptic flow produced in Au-Au collisions at different beam energies. HYDJET++ model shows a good match with the experimentally observed particle ratios, which validates the Cleymans-Reidlich parameterization of $T_{ch}$ and $\mu_{B}$ for lower collision energies. HYDJET++ calculations of $p_{T}$ spectra are in agreement with experimental data for central and semi-central collisions. The lower energies of the BES have a higher chemical potential and a lower freeze-out, these conditions affect the $p_{T}$ distribution of p and $\bar{p}$. A large $\mu_{B}$ favors the thermal proton production while reducing the anti-proton production. The jet contribution also reduces due to a decrease in inelastic cross-section ($\sigma_{NN}^{inel}$) with beam energy. The interplay between these effects can be seen in the $p_{T}$ spectra of $\bar{p}$ where the jet contributions affect the low $p_{T}$ spectra. The model fails to describe the p and $\bar{p}$ spectra in peripheral collisions, this can be attributed to the centrality dependence of $\mu_{B}$, which is observed for these collision energies.
We have re-interpreted elliptic flow ($v_{2}$) based on the scaling of azimuthal spatial anisotropy. This approach allows us to calculate $v_{2}$ without having to track the whole hydrodynamic evolution of the system. The scaling between initial and final spatial anisotropy, referred to as “$k$” in this work, is used to calculate $v_{2}$. The approach can successfully describe the experimental data up to 11.5 GeV. It also preserves the essential aspects of $v_{2}$ such as mass ordering at both low $p_{T}$ ($v_{2}(\pi^{+})>v_{2}(k^{+})>v_{2}(p)$) and high $p_{T}$ ($v_{2}(p)>v_{2}(\bar{p})$). The value of “$k$” is found to be a function of beam energy and can be parameterized as a function of $\ln{\sqrt{s_{NN}}}$. The primary mesons are found to be more sensitive to $k$ than the baryons. A different value of “$k$” is required for strange and non-strange hadrons to better describe the data, which suggests that the relatively heavier strange quarks have smaller spatial and momentum anisotropy. However, this observation is empirical in nature, and more detailed studies are needed to have a better understanding of the transport properties of strange quarks.
@@ Acknowledgements.
Data availabilty
The data will be made available on request.